What is the internal energy of 2.0 mol of an ideal monatomic gas at 300 K?
J
Answer #1

U= Cv *n * R * T
Here
U= Internal Energy
Cv = 3/2 for a monoatomic gas
n= 1 mole
R = 8.314 J/(K * Moles)
T=300K

Therefore, U = 3741.3 Joules
Answer #2

U= ?v *n * R * T = 3/2 * 2*8.314*300 = 7482.6 J

Here
U= Internal Energy
?v = 3/2 for a monoatomic gas
n= 2 mole
R = 8.314 J/(K * Moles)
T=300